# VideoAds for Fast-Paced Video Understanding: Where Opensource Foundation Models Beat GPT-4o & Gemini-1.5 Pro

### v1

@@ 5.1 Impact of Speech Audio Transcript
To evaluate the influence of audio transcripts on model performance, we generate subtitles for each video using OpenAI Whisper [27]. We further test the performance of the model with transcripts and the results are shown in Table 4. Comparing transcript-enhanced models to those using only frame-based information, we observe that introducing audio transcripts significantly improves performance across all MLLMs. This suggests that many MLLMs for video are focusing on visual information, while the important cross-modality reasoning capability remains underexplored.
@@ 5.2 Impact of Chain of Thought
Previous research has shown that Chain of Thought (CoT) can lead to more accurate responses from LLMs. However, the research on the influence of CoT on the multimodality data remains limited, and most studies focus on the image data [6, 37]. Here we present the first study of CoT’s influence on video MLLMs using our challenging dataset. Specifically, we select 100 challenging QA pairs and manually provide detailed reasoning steps for each questions. Next, MLLMs must answer these reasoning questions step by step and provide final answers to the original challenging reasoning questions.
@@
Interestingly, the influence of CoT splits for two kinds of models: Generally, larger models tend to benefit more from the CoT as shown in the NLP field [11, 36]. We observe that LongVA-7B [47], LLava-Next-32B [48], and LLava-Video-72B [48] show a performance drop when dealing long reasoning contexts provided by CoT. However, models like Qwen2.5-VL [34] and GPT-4o [13] show significant performance improvements. Particularly, GPT-4o increases its reasoning accuracy from 61% to 70%, indicating strong reasoning performance given CoT. This also underscores the importance of introducing long context and multi-round VQA in MLLM training [37, 6].
@@ 6 Conclusion
Supplementary Material
I Video filtering criteria
The filtering criteria for our VideoAds during the video selection included: i) Non-advertisement videos (e.g., user-generated content, review videos, or behind-the-scenes footage). ii) Short videos (less than 30 seconds), which typically lack the necessary narrative progression for meaningful temporal reasoning. iii) Low-quality videos (e.g., advertisements consisting solely of several static images, lacking dynamic scene transitions). In particular, steps i and iii filtering are performed by human experts, while step ii filtering is conducted automatically.
II Further discussion on the video complexity
Here we further provide the influence of time duration $d$ on the $V_{cpx}$, and we can observe that increasing time duration $V_{cpx}$ tends to increase due to more variance in the long time unless it already reaches the maximum the video durations. And regardless of the time duration, the VideoAds can significantly out perform other datasets in terms of video complexity.
5 10 15 20 25 30
TGIF 20.79 21.10 21.11 21.11 21.11 21.11
ActivityNet 37.91 40.82 41.74 42.19 42.44 42.60
TVQA 41.85 45.06 46.11 46.64 46.94 47.12
MSVD 25.36 26.55 26.68 26.70 26.70 26.70
MSRVTT 33.25 35.62 36.00 36.06 36.08 36.09
EgoSchema 25.88 31.04 33.61 35.28 36.51 37.47
AutoEval-Video 26.43 30.46 31.54 31.95 32.14 32.26
TempCompass 16.50 20.21 21.23 21.55 21.63 21.64
NExTVideo 19.54 24.01 26.09 27.26 27.97 28.43
Video-Bench 23.10 26.88 28.31 29.08 29.59 29.97
MVBench 17.39 20.43 21.71 22.36 22.67 22.83
VideoMME-S 52.06 56.10 57.44 58.09 58.47 58.72
VideoMME-M 32.98 42.38 47.08 50.05 52.14 53.71
VideoAds 58.71 67.40 70.65 72.32 73.31 73.89
Table 5: The influence of time duration in calculating the video complexity, with the increasing time duration $V_{cpx}$, tends to increase due to more variance in the long term.
It is also worth noticing that the video complexity score only measures the complexity of the video, and can not fully represent the complexity in terms of VQA benchmarks. For example, TempCompass dataset [19] show the challenging of time order in the video understanding for MLLMs. While the video itself can be easy, there exists the possibility that the VQA question can be still challenging.
III Duration Distribution of VideoAds
Here we provide the distribution of video durations in Figure 5, interesting, we can find that most advertisement videos are particularly focusing on 30 seconds, 60 seconds, and 90 seconds. This aligns with the real work advertisement video distribution given the property of advertisement markers.
Figure 5: The video duration distribution of proposed VideoAds. We can find most advertisement videos are particularly focusing on 30 seconds, 60 seconds, and 90 seconds.
IV Failure case study
Here we provide several failure case by common MLLMs study in VideoAds benchmark as shown in Figure 6,7,8,9,12. Interestingly, for some reasoning questions when it is against common sense like Figure 6, MLLMs tend to generate answers based on common sens rather than visual information.
Figure 6: Selected case visualization with MLLMs prediction
Figure 7: Selected case visualization with MLLMs prediction
Figure 8: Selected case visualization with MLLMs prediction
Figure 9: Selected case visualization with MLLMs prediction
Figure 10: Selected case visualization with MLLMs prediction
Figure 11: Selected visualization for case with low complexity with MLLMs prediction
Figure 12: Selected visualization for case with low complexity with MLLMs prediction
V Why not CLIP?
The reason for choosing DINO rather than CLIP [26] is that DINO is more focused on the changes of the low-level visual components, while CLIP is more noisy due to semantic encoding [33]. Here we provide several examples in Figure 13 where the CLIP provides one high video complexity but the video actually contains low complexity.
Figure 13: Some examples that CLIPs generate over-estimation for the video complexity.
VI More discussion in related work
Traditional video evaluation benchmarks focus on one specific task by collecting data from the corresponding domain. For example, ActivityNet-QA [44] focuses on human activity recognition, and MSVD-QA [41] are generated by video description datasets and focus on action and object recognition. Along with the rapid progress of the current LLMs [3, 10], the recent VLM enables the process of complex video in more than one single motion or short video clip [9]. Correspondingly, the recent video evaluation benchmarks are more focused on more challenging and comprehensive video understanding. Video-MME [8] is one widely used comprehensive video evaluation benchmark containing different domains and various video lengths (from seconds to hours). MVBench [17] covers 20 video tasks for spatial understanding and temporal understanding. This benchmark also builds a semiautomatic reannotation pipeline using ChatGPT for existing video datasets with original annotations. MMBench-Video [18] focuses on free-form questions from lengthy YouTube videos and introduces GPT-4 [1] for automated assessment. Video-Bench [22] includes three types of questions, including video-exclusive Understanding, prior Knowledge-based question-answering, and comprehension and decision-making based on the real-world video. TempCompass [19] focuses on on the temporal perception ability of MLLM by collecting videos that share the same static content but differ in a specific temporal aspect. AutoEval-Video [5] develop a novel adversarial annotation mechanism and constructs open-ended video-questions across 9 skill dimensions. Another interesting recent work named Video-MMMU [12] focuses on long professional videos that systematically evaluate knowledge acquisition capabilities.

### v2

@@ 5.1 Impact of Speech Audio Transcript
To evaluate the influence of audio transcripts on model performance, we generate subtitles for each video using OpenAI Whisper [26]. We further test the performance of the models with transcripts and the results are shown in Table 4. Comparing transcript-enhanced models to those using only frame-based information, we observe that introducing audio transcripts significantly improves performance across all MLLMs. This suggests that many MLLMs for video are focusing on visual information, while the important cross-modality reasoning capability remains underexplored.
@@ 5.2 Impact of Chain of Thought
The research on the influence of CoT on the multimodality data remains limited, and most studies focus on the image data [4, 36]. Here we present the first study of CoT’s influence on video MLLMs using our challenging dataset. In our CoT prompting, we provided explicit intermediate questions to guide reasoning, rather than using generic ’think step-by-step’ instructions. For example, for a visual reasoning task, we might pose intermediate questions like, ’Q1: What product appears in the first 5 seconds?’ followed by ’Q2: What emotion does the character express at the end?’ The model’s answers to these intermediate questions were then used as context to generate the final response to the main question. We adopted this guided setup because many current MLLMs do not reliably support open-ended, multi-round conversational reasoning from a single generic prompt.
@@
Interestingly, the influence of CoT differs across models: Generally, larger models tend to benefit more from the CoT as shown in the NLP field [9, 35]. We observe that LongVA-7B [47], LLava-Next-32B [48], and LLava-Video-72B [48] show a performance drop when dealing long reasoning contexts provided by CoT. However, models like Qwen2.5-VL [33] and GPT-4o [11] show significant performance improvements. Particularly, GPT-4o increases its reasoning accuracy from 61% to 70%, indicating strong reasoning performance given CoT. This also underscores the importance of introducing long context and multi-round VQA in MLLM training [36, 4].